# Large Language Models for Information Retrieval: A Survey

### v4

@@ Abstract
Large Language Models; Information Retrieval; Query Rewriter; Reranking; Reader; Fine-tuning; Prompting
@@ 3.1 Rewriting Scenarios
By leveraging the capabilities of LLMs, researchers have been able to generate a variety of formats for rewritten queries, such as questions [78, 80, 94, 91, 88, 89, 90], answer-incorporated passages [65, 77, 83, 84, 85, 75, 76, 87, 95], and keywords [82, 96, 81, 80]. Comprehensive details for each format are available in the following part.
@@ 3.3.1 Prompting
$\bullet$ Chain-of-thought prompting. CoT prompting [97] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [78, 76]. In conversational search, the process of query rewriting is multi-turn, which means queries should be refined step-by-step with the interaction between search engines and users. This process naturally coincides with the CoT process. As shown in Figure 3, users can conduct the CoT process by adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one round in query rewriting, so CoT could only be accomplished in a simple and coarse way. For example, as shown in Table II, researchers add “Give the rationale before answering” in the instructions to prompt LLMs to think deeply [76].
@@ 3.3.3 Reinforcement Learning
Query rewriters typically serve as intermediaries for retrieval systems, implying that they do not operate independently, and there is no specific loss attributed solely to the query rewriting process. In such a context, Reinforcement Learning (RL) presents an alternative training mechanism for the query rewriter. Query rewriters can receive feedback signals from ranking models [88] or downstream LLM readers [80]. For instance, ranking scores can be utilized to construct good-bad pairs for DPO [101] training. Ma et al. [80] proposed a method that first generates answers from LLMs and then uses the QA evaluation results as training signals. In a similar manner, BEQUE [81] introduced an offline feedback system capable of providing a quality score for each query based on the set of products it retrieved. Such an application of Reinforcement Learning mechanisms aligns the objective of query rewriters more closely with the goals of downstream tasks, thereby enhancing the overall performance of the system.
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. According to the usage of LLMs, existing LLM-based reranking methods can be divided into three paradigms: utilizing LLMs as supervised rerankers, utilizing LLMs as unsupervised rerankers, and utilizing LLMs for training data augmentation. These paradigms are summarized in Table V and will be introduced in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [99]).
@@
Supervised
Rerankers
Enc-only monoBERT [12]
Enc-dec
monoT5[13], Ju et al. [142], DuoT5 [143],
RankT5 [144], ListT5 [145]
Dec-only
RankLLaMA [122], TSARankLLM [146],
Q-PEFT [147],  Zhang et al. [148],
PE-Rank [149]
Unsupervised
Rerankers
Pointwise
Liang et al. [150], Zhuang et al. [151],
Guo et al. [152], Sachan et al. [153],
Zhuang et al. [154], Sun et al. [155],
Co-Prompt [156], DemoRank [157],
PARADE [158]
Listwise
RankGPT [159],  Ma et al. [160],
Tang et al. [161], TourRank [162],
Parry et al. [163]
Pairwise
PRP [164], Zhuang et al. [165],
PRP-Graph [166], Yan et al. [167]
Data
Augmentation
ExaRanker [168], ExaRanker-Open [169],
InPars-Light [170], Askari et al. [171],
Askari et al. [172], RankVicuna [173],
RankZephyr [174], Sun et al. [175]
@@ 5.1.1 Encoder-only
The encoder-based rerankers represent a significant turning point in applying LLMs to document ranking tasks. They demonstrate how some pre-trained language models (e.g., BERT [55]) can be fine-tuned to deliver highly accurate relevance predictions. A representative approach is monoBERT [12], which transforms a query-document pair into a sequence “[CLS] query [SEP] document [SEP]” as the model input and calculates the relevance score by feeding the “[CLS]” representation into a linear layer. The reranking model is optimized based on the cross-entropy loss.
@@ 5.1.2 Encoder-Decoder
In this field, existing studies mainly formulate document ranking as a generation task and optimize an encoder-decoder-based reranking model [13, 142, 143, 144]. Specifically, given the query and the document, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”. During inference, the query-document relevance score is determined based on the logit of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is calculated as the probability of the “true” token. The following method [142] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [143] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
@@ 5.1.3 Decoder-only
Recently, there have been some attempts [122, 146, 147, 148, 149] to rerank documents by fine-tuning decoder-only models (such as LLaMA). For example, RankLLaMA [122] proposes formatting the query-document pair into a prompt “query: {query} document: {document} [EOS]” and utilizes the last token representation for relevance calculation. Besides, TSARankLLM [146] has been proposed to bridge the gap between LLMs’ conventional training objectives and the specific needs of document ranking through two-stage training. The first stage involves continuously pretraining LLMs using a large number of relevant text pairs collected from web resources, helping the LLMs to naturally generate queries relevant to the input document. The second stage focuses on improving the model’s text ranking performance using high-quality supervised data and well-designed loss functions.  Peng et al. [147] propose a query-dependent parameter efficient fine-tuning (Q-PEFT) approach for ranking, which helps the LLM generate true queries based on given documents. Different from these pointwise rerankers [122, 146, 147],  Zhang et al. [148] and  Liu et al. [149] proposes to train a listwise reranker that directly outputs a reranked document list. Specifically,  Zhang et al. [148] first demonstrate that existing pointwise datasets (such as MS MARCO [99]), which only contain binary query-document labels, are insufficient for training efficient listwise rerankers. Then, they propose to use the ranking results of existing ranking systems (such as Cohere rerank API) as gold rankings to train a listwise reranker based on Code-LLaMA-Instruct.  Liu et al. [149] propose PE-Rank, which compresses each document in the list into a single embedding and then inputs these document embeddings into reranker, which significantly reduces the input length and improves the efficiency of reranker.
Figure 5: Three types of unsupervised reranking methods: (a) pointwise methods that consist of relevance generation (upper) and query generation (lower), (b) listwise methods, and (c) pairwise methods.
@@ 5.2.1 Pointwise methods
TABLE VI: The comparison between different methods. NN denotes the number of documents to rerank. The Complexity, Logits, and Batch represent the computational complexity, whether accesses LLM’s logits, and whether allows batch inference respectively. $k$ is the constant in sliding windows strategy. As for the Performance, we use NDCG@10 as a metric, and the results are calculated by reranking the top 100 documents retrieved by BM25 on TREC-DL2019 and TREC-DL2020. The best model is in bold while the second-best is marked with an underline. The results come from previous study [164]. *Since the parameters of ChatGPT have not been released, its model parameters are based on public estimates [177].
Methods LLM Size Properties Performance
Complexity Logits Batching TREC-DL19 TREC-DL20
@@
Unsupervised-Pointwise Query Generation [153] FLAN-UL2 20B $O(N)$ $\checkmark$ $\checkmark$ 58.95 60.02
@@
Unsupervised-Listwise RankGPT3.5 [159] gpt-3.5-turbo 154B* O(kN)O(k*N) 65.80 62.91
@@
Unsupervised-Pairwise PRP-Allpair [164] FLAN-UL2 20B $O(N^{2})$ $\checkmark$ $\checkmark$ 72.42 70.68
@@ 5.2.2 Listwise Methods
Although listwise methods have yielded promising performance, they still suffer from some weaknesses: (1) The performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25 [159], revealing positional bias issues in the listwise ranking of LLMs. (2) The use of sliding windows limits the number of documents that can be ranked each time, and the dependency between adjacent windows prevents parallelization of LLM inference, thereby reducing the efficiency of reranking. Recently, some studies have attempted to mitigate these issues. Tang et al. [161] introduce a permutation self-consistency method, which involves shuffling the list in the prompt and aggregating the generated results to achieve a more accurate and a positionally unbiased ranking.  Chen et al. [162] introduce a tournament mechanism into listwise ranking and propose TourRank, which parallelizes the reranking process through intelligent grouping and use a tournament-like points system to reduce the impact of the initial document order.  Parry et al. [163] propose a parallelizable partitioning algorithm for listwise ranking, which also aims at mitigating efficiency issues.  Reddy et al. [178] propose a novel listwise reranking approach which leverages the output logits of the first generated identifier to derive a ranked passage list.
@@ 5.2.3 Pairwise Methods
In pairwise methods, LLMs are given a prompt that consists of a query and a document pair (see Figure 5 (c)). Then, they are instructed to generate the identifier of the document with higher relevance. To rerank all candidate documents, aggregation methods like AllPairs are used [164]. AllPairs first generates all possible document pairs, yields discrete judgments for each pair (e.g., Document 1 or Document 2), and aggregates a final relevance score for each document. Efficient sorting algorithms, such as heap sort and bubble sort, are employed to speed up the ranking process. These sorting algorithms utilize efficient data structures to compare document pairs selectively and elevate the most relevant documents to the top of the ranking list, which is particularly useful in top-$k$ ranking. Experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than those typically employed in listwise methods (e.g., GPT3.5). Building on the pairwise prompting approach,  Luo et al. [166] introduce an innovative scoring unit that leverages the generation probability of judgments instead of discrete judgments, and further design a graph-based aggregation approach to obtain a final relevance score for each document. Besides, the pairwise comparison can also be utilized as a post-processing step to adjust the relevance scores generated by the pointwise LLM reranker [167].
@@ 5.3 Utilizing LLMs for Training Data Augmentation
Recently, many studies [173, 174, 175] have also attempted to distill the document ranking capability of LLMs into a specialized model. RankVicuna [173] proposes to use the ranking list of $\text{RankGPT}_{3.5}$ [159] as the gold list to train a 7B parameter Vicuna model. RankZephyr [174] introduces a two-stage training strategy for distillation: initially applying the RankVicuna recipe to train $\text{Zephyr}{\gamma}$ in the first stage, and then further finetuning it in the second stage with the ranking results from $\text{RankGPT}_{4}$. These two studies not only demonstrate competitive results but also alleviate the issue of ranking results non-reproducibility of black-box LLMs. Besides, researchers [175] have also tried to distill the ranking ability of a pairwise ranker, which is computationally demanding, into a simpler but more efficient pointwise ranker.

### v5

@@ Abstract
Large Language Models; Information Retrieval; Query Rewriter; Reranking; Reader; Fine-tuning; Prompting; Agent
@@ 2.2 Large Language Models
Recently, research has focused on enhancing LLM capabilities by improving their reasoning and inference-time processes. Large reasoning models (LRMs) are an evolution of LLMs specifically designed to excel at complex logical tasks such as mathematics and coding. These models often employ advanced training methodologies, including reinforcement learning, to develop sophisticated self-reflection and planning mechanisms. They are designed to generate a detailed “thinking process” before providing a final answer, which has shown improved performance on reasoning benchmarks. This focus on improving the reasoning process is highly relevant to test-time scaling, a paradigm that allocates additional computational resources during inference to improve model performance without increasing model size during pre-training. This can involve generating multiple outputs in parallel and selecting the best one, or employing sequential methods where the model iteratively refines its own output. By achieving promising results on expert-level human challenges, these LRMs provide a new paradigm for IR systems.
@@ 3.1 Rewriting Scenarios
Beyond traditional retrieval scenarios, query rewriting is also widely used in a variety of practical domains. In the context of agents, effectively identifying the most relevant tools for a given task becomes a key bottleneck as the toolset size grows, hindering reliable tool utilization. To address this, current study [94] propose generating a diverse set of synthetic queries that comprehensively cover different aspects of the query space associated with each tool document during the tool indexing phase. On the other hand, in clinical terminology normalization, recent study [95] also leverage large language models to decompose and reconstruct complex diagnostic mentions, improving the mapping to standard terms through a ”retrieve-and-rank” framework to enhance overall performance.
By leveraging the capabilities of LLMs, researchers have been able to generate a variety of formats for rewritten queries, such as questions [78, 80, 96, 91, 88, 89, 90, 97], answer-incorporated passages [65, 77, 83, 84, 85, 75, 76, 87, 98], and keywords [82, 99, 81, 80]. Comprehensive details for each format are available in the following part.
@@ 3.3.1 Prompting
$\bullet$ Chain-of-thought prompting. CoT prompting [100] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [78, 76, 104, 105]. In conversational search, the process of query rewriting is multi-turn, which means queries should be refined step-by-step with the interaction between search engines and users. This process naturally coincides with the CoT process. As shown in Figure 3, users can conduct the CoT process by adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one round in query rewriting, so CoT could only be accomplished in a simple and coarse way. For example, as shown in Table II, researchers add “Give the rationale before answering” in the instructions to prompt LLMs to think deeply [76].
@@ 3.3.3 Reinforcement Learning
Query rewriters typically serve as intermediaries for retrieval systems, and as such, they lack a dedicated or independent loss function for optimization. In this context, reinforcement learning (RL) presents an alternative training paradigm. The query rewriter can receive feedback signals from donwstream components, such as ranking models [88] or LLM readers [80]. For instance, ranking scores can be utilized to construct good-bad pairs for direct preference optimization [106] training. Similarly, Ma et al. [80] propose to generate answers from LLMs and then uses the results of a QA evaluation as training signals. Another approach, BEQUE [81], introduces an offline feedback system that assigns a quality score to each query based on the set of products it retrieves.
Recently, inspired by the rule-based reward system proposed by DeepSeek-R1 [107], some studies have explored using retrieval metrics as the reward to optimize query generators. For example, DeepRetrieval [108] proposes a RL approach that trains LLMs for query generation by using retrieval metrics as rewards without the need for supervised data. This method reinforces the query generator to produce queries that maximize retrieval performance. These RL mechanisms align the objective of query rewriters more closely with the goals of downstream tasks, thereby enhancing the overall performance of the system.
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. According to the usage of LLMs, existing LLM-based reranking methods can be divided into four paradigms: utilizing LLMs as supervised rerankers, utilizing LLMs as unsupervised rerankers, utilizing LLMs for training data augmentation and reasoning-intensive rerankers. These paradigms are summarized in Table V and will be introduced in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [102]).
@@
Supervised Rerankers Enc-only monoBERT [12]
Enc-dec monoT5 [13], Ju et al. [149], DuoT5 [150], RankT5 [151], ListT5 [152]
Dec-only RankLLaMA [129], TSARankLLM [153], Q-PEFT [154], Zhang et al. [155], PE-Rank [156]
Unsupervised Rerankers Pointwise Liang et al. [157], Zhuang et al. [158], Guo et al. [159], Sachan et al. [160], Zhuang et al. [161], Sun et al. [162], Co-Prompt [163], DemoRank [164], PARADE [165]
Listwise RankGPT [166], Liu et al. [167], CoRanking [168], Ma et al. [169], Tang et al. [170], TourRank [171], Parry et al. [172], APEER [173]
Pairwise PRP [174], Zhuang et al. [175], PRP-Graph [176], Yan et al. [177]
Data Augmentation ExaRanker [178], ExaRanker-Open [179], InPars-Light [180], Askari et al. [181], Askari et al. [182], RankVicuna [183], RankZephyr [184], Sun et al. [185]
Reasoning-intensive Rerankers ReasonRank [186], Rank1 [187], Rank-K [188], Rearank [189], Rank-R1 [190], TFRank [191]
@@ 5.1.1 Encoder-only
The encoder-based rerankers represent a significant turning point in applying LLMs to document reranking tasks. They demonstrate how some pre-trained language models (e.g., BERT [55]) can be fine-tuned to deliver highly accurate relevance predictions. A representative approach is monoBERT [12], which transforms a query-document pair into a sequence “[CLS] query [SEP] document [SEP]” as the model input and calculates the relevance score by feeding the “[CLS]” representation into a linear layer. The reranking model is optimized based on the cross-entropy loss.
@@ 5.1.2 Encoder-Decoder
In this field, existing studies mainly formulate document reranking as a generation task and optimize an encoder-decoder-based reranking model [13, 149, 150, 151]. Specifically, given the query and the document, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”. During inference, the query-document relevance score is determined based on the logit of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is calculated as the probability of the “true” token. The following method [149] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [150] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
@@ 5.1.3 Decoder-only
Recently, there have been some attempts [129, 153, 154, 155, 156] to rerank documents by fine-tuning decoder-only models (such as LLaMA). For example, RankLLaMA [129] proposes formatting the query-document pair into a prompt “query: {query} document: {document} [EOS]” and utilizes the last token representation for relevance calculation. Besides, TSARankLLM [153] has been proposed to bridge the gap between LLMs’ conventional training objectives and the specific needs of document reranking through two-stage training. The first stage involves continuously pretraining LLMs using a large number of relevant text pairs collected from web resources, helping the LLMs to naturally generate queries relevant to the input document. The second stage focuses on improving the model’s text ranking performance using high-quality supervised data and well-designed loss functions.  Peng et al. [154] propose a query-dependent parameter efficient fine-tuning (Q-PEFT) approach for ranking, which helps the LLM generate true queries based on given documents. Different from these pointwise rerankers [129, 153, 154],  Zhang et al. [155] and  Liu et al. [156] proposes to train a listwise reranker that directly outputs a reranked document list. Specifically,  Zhang et al. [155] first demonstrate that existing pointwise datasets (such as MS MARCO [102]), which only contain binary query-document labels, are insufficient for training efficient listwise rerankers. Then, they propose to use the ranking results of existing ranking systems (such as Cohere rerank API) as gold rankings to train a listwise reranker based on Code-LLaMA-Instruct.  Liu et al. [156] propose PE-Rank, which compresses each document in the list into a single embedding and then inputs these document embeddings into reranker, which significantly reduces the input length and improves the efficiency of reranker.
Figure 5: Three types of unsupervised reranking methods: (a) pointwise methods that consist of relevance generation (upper) and query generation (lower), (b) listwise methods, and (c) pairwise methods. The “Demonstrations” represents the few-shot demonstrations whose format are same as the current input.
@@ 5.2.1 Pointwise methods
TABLE VI: The comparison between different LLM-based reranking methods. NN denotes the number of documents to rerank. The “complexity”, “logits”, and “batch” represent the computational complexity, whether accesses LLM’s logits, and whether allows batch inference respectively. $k$ is the constant in sliding windows strategy. As for the performance, we use NDCG@10 as a metric, and the results are calculated by reranking the top-100 documents retrieved by BM25 on TREC-DL2019 and TREC-DL2020. The best model is in bold while the second-best is underlined. The results come from previous study [174]. *Since the parameters of ChatGPT have not been released, its model parameters are based on public estimates [193].
Method LLM Size Properties Performance
Complexity Logits Batching TREC-DL19 -DL20
@@
Unsupervised Pointwise Query Generation [160] FLAN-UL2 20B $O(N)$ $\checkmark$ $\checkmark$ 58.95 60.02
@@
Unsupervised Listwise RankGPT3.5 [166] ChatGPT 154B* O(kN)O(k*N) 65.80 62.91
@@
Unsupervised Pairwise PRP-Allpair [174] FLAN-UL2 20B $O(N^{2})$ $\checkmark$ $\checkmark$ 72.42 70.68
@@ 5.2.2 Listwise Methods
Although listwise methods have yielded promising performance, they still suffer from some weaknesses: (1) The performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25 [166], revealing positional bias issues in the listwise ranking of LLMs. (2) The use of sliding windows limits the number of documents that can be ranked each time, and the dependency between adjacent windows prevents parallelization of LLM inference, thereby reducing the efficiency of reranking. Recently, some studies have attempted to mitigate these issues. Tang et al. [170] introduce a permutation self-consistency method, which involves shuffling the list in the prompt and aggregating the generated results to achieve a more accurate and a positionally unbiased ranking.  Chen et al. [171] introduce a tournament mechanism into listwise ranking and propose TourRank, which parallelizes the reranking process through intelligent grouping and use a tournament-like points system to reduce the impact of the initial document order.  Parry et al. [172] propose a parallelizable partitioning algorithm for listwise ranking, which also aims at mitigating efficiency issues.  Reddy et al. [194] propose a novel listwise reranking approach which leverages the output logits of the first generated identifier to accelerating reranking process. To optimize the listwise reranking prompt,  Jin et al. [173] propose a novel automatic prompt engineering algorithm APEER, which generates prompts through feedback and preference optimization.  Liu et al. [167] comprehensively discuss the benefits of long-context LLMs for listwise ranking and introduce a novel full reranker which performs better than the sliding window reranker while also being more efficient and having lower API cost.  Liu et al. [168] propose a collaborative ranking framework CoRanking which combines small and large listwise rerankers for more efficient and effective passage ranking. They also design a novel passage order adjuster to mitigate the sensitivity of listwise reranker to the input document order.
@@ 5.2.3 Pairwise Methods
In pairwise methods, LLMs are given a prompt that consists of a query and a document pair (see Figure 5 (c)). Then, they are instructed to generate the identifier of the document with higher relevance. To rerank all candidate documents, aggregation methods like AllPairs are used [174]. AllPairs first generates all possible document pairs, yields discrete judgments for each pair (e.g., Document 1 or Document 2), and aggregates a final relevance score for each document. Efficient sorting algorithms, such as heap sort and bubble sort, are employed to speed up the ranking process. These sorting algorithms utilize efficient data structures to compare document pairs selectively and elevate the most relevant documents to the top of the ranking list, which is particularly useful in top-$k$ ranking. Experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than those typically employed in listwise methods (e.g., GPT3.5). Building on the pairwise prompting approach, several ranking method variants have been proposed.  Luo et al. [176] introduce an innovative scoring unit that leverages the generation probability of judgments instead of discrete judgments, and further design a graph-based aggregation approach to obtain a final relevance score for each document.  Sinhababu et al. [195] and propose to utilize few-shot in-context demonstrations to improve the performance of pairwise ranking.  Yan et al. [177] utilize the pairwise comparison as a post-processing step to adjust the relevance scores generated by the pointwise LLM reranker.
@@ 5.3 Utilizing LLMs for Training Data Augmentation
Recently, many studies [183, 184, 185] have also attempted to distill the document reranking capability of LLMs into a specialized model. RankVicuna [183] proposes to use the ranking list of $\text{RankGPT}_{3.5}$ [166] as the gold list to train a 7B parameter Vicuna model. RankZephyr [184] introduces a two-stage training strategy for distillation: initially applying the RankVicuna recipe to train $\text{Zephyr}{\gamma}$ in the first stage, and then further finetuning it in the second stage with the ranking results from $\text{RankGPT}_{4}$. These two studies not only demonstrate competitive results but also alleviate the issue of ranking results non-reproducibility of black-box LLMs. Besides, researchers [185] have also tried to distill the ranking ability of a pairwise ranker, which is computationally demanding, into a simpler but more efficient pointwise ranker.
5.4 Reasoning-intensive Rerankers
Recent breakthroughs in Large Reasoning Models (LRMs) like DeepSeek-R1 [107] have demonstrated exceptional capabilities across many NLP tasks. These models significantly improve the answer accuracy in many complex NLP tasks (e.g., math and coding) through explicit step-by-step reasoning chains during inference. This capability holds particular promise for document reranking, where precise understanding of query intent and cross-document comparison are critical for relevance assessment.
Motivated by these advancements, emerging research has explored injecting reasoning ability into document rerankers. For example,  Weller et al. [187] and  Yang et al. [188] propose to apply DeepSeek-R1 as a teacher model to distill its reasoning process into smaller rerankers.  Zhang et al. [189] and  Zhuang et al. [190] propose to use reinforcement learning algorithm to optimize reranker based on rule-based reward. TFRank [191] introduces a “think-free” pointwise ranker that leverages reasoning during training while eliminating intermediate reasoning steps at inference, significantly improving the reasoning efficiency. While effective, these rerankers are primarily trained on traditional web search data MSMARCO, making them difficult to generalize to many complex and reasoning-intensive ranking benchmarks [196]. To address the scarcity of reasoning-intensive training data, Liu et al. [186] propose an automated data synthesis framework and generate 13K high-quality reasoning-intensive training data covering diverse search scenarios. They further propose a two-stage “SFT+RL” training framework, to empower LLM with strong reasoning and ranking abilities. Their ReasonRank model has achieved state-of-the-art performance on many reasoning-intensive IR benchmarks such as BRIGHT [196] and R2MED [197].